OpenStudy (anonymous):

Evaluate 2x^2 − 5y ---------- 3x − y when x = 2 and y = -4.

OpenStudy (anonymous):

14/5 or 2.8

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\(\huge \frac{2x^2 − 5y}{ 3x − y }=\frac{2\times2^2-5\times(-4)}{3\times 2-(-4)}=\frac{8+20}{6+4}=\frac{28}{10}=\frac{14}{5} \)

OpenStudy (anonymous):

no prob